OpenStudy (anonymous):

Can someone explain to me how to answer this problem in simple ways? (I get really confused when it comes to math) A communications satellite transmits a radio wave at a frequency of 5 • 109 Hz. What is the signal’s wavelength? Assume the wave travels in a vacuum.

OpenStudy (phi):

this is physics \[ f\lambda = c \] where f is the frequency, lambda is the wavelength, and c is the speed of light (in a vacuum)

OpenStudy (phi):

so divide f into c to find lambda.

OpenStudy (apoorvk):

All electromagnetic waves have the same speed 'c' which equals 3x10^8 m/sec and radio waves are definitely electromagnetic waves. Now, wave speed = wavelength * frequency You know the wave speed and the frequency, can you find out the wavelength?

OpenStudy (anonymous):

Yes! thank you very much both of you!

OpenStudy (phi):

\[ \frac{3\cdot 10^8 \text{m/s} }{5\cdot 10^9 \text{ 1/s}} = \frac{3}{5}\cdot 10^{-1} \text{ m} = 0.6\cdot 10^{-1} m= 0.06 m\]

OpenStudy (apoorvk):

You're welcome!